# Optimal Power Flow Pursuit via Feedback-based Safe Gradient Flow

### v2

@@ Abstract
This paper considers the problem of controlling inverter-interfaced distributed energy resources (DERs) in a distribution grid, to achieve operational and performance goals with limited system-level information. We develop an online feedback optimization method to drive the DERs’ power setpoints to solutions of an AC optimal power flow (OPF) problem based only on voltage measurements (and without requiring measurements of the power consumption of non-controllable assets). The proposed method – grounded on the theory of control barrier functions – is based on a continuous approximation of the projected gradient flow, appropriately modified to accommodate measurements from the power network. We provide results in terms of local exponential stability, and assess the robustness to errors in the measurements and in the system Jacobian matrix. We show that the proposed method ensures anytime satisfaction of the voltage constraints when no model and measurement errors are present; if these errors are present and are small, the voltage violation is practically negligible. We also discuss extensions of the framework to virtual power plant setups. Numerical experiments on a 93-bus distribution system and with realistic load and production profiles show a superior performance in terms of voltage regulation relative to existing methods.
@@ I Introduction
Traditional techniques based on solving an AC optimal power flow (OPF) problem [3, 4] require collecting information of all non-controllable powers and running an iterative method; this process may be long compared to the fast changing conditions of a modern distribution system [1, 2]; existing Volt/Var techniques may not fully resolve voltage regulation and may in fact increase line currents; recent works on emulating OPF solutions via neural networks can alleviate the computational burden [5, 6], but still require measurements of all the non-controllable powers (which are the inputs to the neural network), and may not even produce feasible power setpoints. In this work, we focus on real-time AC OPF methods [7, 8, 9, 10, 11, 12], and seek new strategies that exhibit strong performance in terms of achieved operational cost and voltage limit satisfaction (both from analytical and numerical standpoints), while using limited system information and measurements. In particular, we seek methods that do not require a complete AC model and knowledge of all the non-controllable powers throughout the nodes of the system.
@@
Contributions. Our contributions are as follows.
@@
(c2) From a theoretical standpoint, we show that the proposed feedback-based SGF renders isolated optimal solutions of the AC OPF problem locally exponentially stable and ensures the anytime satisfaction of the voltage constraints. Moreover, we provide practical exponential stability results when voltage measurements are affected by errors and when the Jacobian matrix of the AC power flow equations is computed only approximately. In this case, if errors are small, the voltage violation is practically negligible.
(c3) We perform numerical experiments on a 93-bus distribution system [24] and with realistic load and solar production profiles from the Open Power System Data. We show that our method shows far superior performance in terms of voltage regulation relative to existing online primal-dual methods and Volt/Var strategies.
@@ II-A Distribution system model
[…]
@@
where $\boldsymbol{s}=\boldsymbol{p}_{\text{net}}+\text{j}\boldsymbol{q}_{\text{net}}\in\mathbb{C}^{N}$, with $\boldsymbol{p}_{\text{net}}$ and $\boldsymbol{q}_{\text{net}}$ vectors collecting the net active and reactive power injections at nodes $\mathcal{N}$. Note that $\boldsymbol{p}_{\text{net}}$ and $\boldsymbol{q}_{\text{net}}$ account for both the powers (injected or consumed) of the DERs and the aggregate powers of the non-controllable loads that are connected to each of the nodes $\mathcal{N}$. In particular, let $\boldsymbol{p}_{l}:=[p_{l,1},\dots,p_{l,N}]^{\top}\in\mathcal{W}_{p},\boldsymbol{q}_{l}:=[q_{l,1},\dots,q_{l,N}]^{\top}\in\mathcal{W}_{q}$ with a compact sets $\mathcal{W}_{p}\subset\mathbb{R}^{N}$ and $\mathcal{W}_{q}\subset\mathbb{R}^{N}$, be vectors collecting the net active and reactive power consumed at the nodes by non-controllable devices (positive when the power is consumed). For the $G$ DERs, consider the vector $\boldsymbol{u}=[p_{1},p_{2},...,p_{G},q_{1},q_{2},...,q_{G}]^{\top}$ collecting their active and reactive powers (with a positive sign denoting generation). Moreover, let ${\mathcal{G}}:=\{1,...,G\}$ be the index for the DERs, and define a function $m:{\mathcal{G}}\rightarrow\mathcal{N}$ which maps a DER index to the node where the DER is connected to. With this notation, note that ${\mathcal{G}}_{n}:=\{i\in{\mathcal{G}}:n=m(i)\}$ is the set of DERs connected at node $n\in\mathcal{N}$. Then, the net active and reactive powers are given by $p_{\text{net},n}=\sum_{j\in{\mathcal{G}}_{n}}p_{j}-p_{l,n}$ and $q_{\text{net},n}=\sum_{j\in{\mathcal{G}}_{n}}q_{j}-q_{l,n}$ at each $n$. In what follows, we consider a set of nodes $\mathcal{M}\subseteq\mathcal{N}$ with cardinality $M=|\mathcal{M}|$ where voltages are to be regulated (if the operator would like to monitor and regulate all the voltages, then $\mathcal{M}=\mathcal{N}$). Finally, let $(p_{i},q_{i})\in\mathcal{U}_{i}$ be the power setpoint for the $i$th DER, where $\mathcal{U}_{i}\subset\mathbb{R}^{2}$ is a compact set of admissible power setpoints. We define $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}\times\ldots\times\mathcal{U}_{G}$ so that $\boldsymbol{u}\in\mathcal{U}$.
@@
There exists a unique continuously differentiable function $H:\mathcal{U}\times\mathcal{W}_{p}\times\mathcal{W}_{q}\rightarrow\mathbb{R}^{M}$ such that, $H_{i}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})=\nu_{i}=|v_{i}|$, for $i\in\mathcal{M}$. The Jacobian $J_{h}(\boldsymbol{u}):=\frac{\partial H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})}{\partial\boldsymbol{u}}$ is locally Lipschitz continuous. $\Box$
@@
Remark II.1 (Model and notation).
@@
II-B OPF for distribution systems
@@
To outline our framework, we start with the following formulation of the OPF problem (we present some extensions later in Section III-C):
@@
i) Strict complementarity condition [29] and the linear independence constraint qualification (LICQ) hold at $\boldsymbol{u}^{*}$.
@@
Problem (4) can be solved using traditional optimization methods for nonlinear programs. However, (c1) these batch methods require collecting measurements of all the non-controllable powers $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$; moreover, (c2) the time required to collect the measurements of the non-controllable powers and run an iterative method to convergence may be long compared to the fast changing conditions of a modern distribution system [1, 2]. Although recent work on neural networks for AC OPF can alleviate the computational burden (see, e.g., the representative works [5, 6]), they still require measurements of all the non-controllable powers $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$ as in (c1), and they often rely on heuristics to return a feasible solution. Motivated by the challenges (c1)(c2) and by the need to generate power setpoints that ensure satisfaction of voltage limits even under uncertain and time-varying operational setups, in this paper we seek to solve to the following problem.
@@
Design an online feedback optimization method that drives the DERs’ power setpoints $\boldsymbol{u}$ to solutions of the problem (4), while ensuring that voltage constraints are always met. The feedback optimization method should use measurements of the voltages instead of requiring knowledge of the non-controllable powers $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$. $\Box$
@@ III-A Feedback-based online algorithm
$F_{\beta}(\boldsymbol{u},\tilde{\boldsymbol{\nu}}):=\arg\min_{\boldsymbol{\theta}\in\mathbb{R}^{2G}}\|\boldsymbol{\theta}+\nabla C_{p}(\boldsymbol{u})+J_{H}(\boldsymbol{u})^{\top}\nabla C_{v}(\tilde{\boldsymbol{\nu}})\|^{2}$
@@
$\hskip 78.24507pt\nabla H_{i}(\boldsymbol{u})^{\top}\boldsymbol{\theta}\leq-\beta\left(\tilde{\nu}_{i}-\bar{V}\right)\,\,\forall i\in\mathcal{M}$
$\hskip 78.24507ptJ_{\ell_{i}}(\boldsymbol{u})^{\top}\boldsymbol{\theta}\leq-\beta\ell_{i}(p_{i},q_{i})\quad\,\,\,\,\forall i\in\mathcal{G}$
where $\tilde{\nu}_{i}$ is a measurement of $|v_{i}|$ at node $i$, $J_{H}(\boldsymbol{u})$ is the Jacobian matrix of $H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$, $\nabla H_{i}(\boldsymbol{u})=[\{(J_{H}(\boldsymbol{u}))_{i,j}\}_{j\in\mathcal{G}}]^{\top}$ is a $2G\times 1$ vector collecting the entries of $J_{H}(\boldsymbol{u})$ in the $i$th row and columns corresponding to nodes in $\mathcal{G}$, $J_{\ell_{i}}(p_{i},q_{i})$ is the Jacobian of $(p_{i},q_{i})\mapsto\ell_{i}(p_{i},q_{i})$, $\beta>0$ is a design parameter, and $\eta>0$ is the controller gain. For given $\boldsymbol{u}$ and $\tilde{\boldsymbol{\nu}}$, problem (8) is a convex quadratic program (QP) with a strongly convex cost; it can be efficiently solved using standard or high-performance embedded solvers for QPs, e.g., [31].
The online feedback optimization algorithm (7) is inspired by CBFs methods [22] and the safe gradient flow in [23]; we provide more details on the CBF-based design in Appendix A. In particular, (7) is an approximation of the projected gradient flow $\dot{\boldsymbol{u}}=\text{proj}_{T_{\mathcal{F}}(\boldsymbol{u})}\{-\nabla C_{p}(\boldsymbol{u})-J_{H}^{\top}\nabla C_{v}(H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}))\}$, where $T_{\mathcal{F}}(\boldsymbol{u})$ is the tangent cone of $\mathcal{F}(\boldsymbol{u})$ at $\boldsymbol{u}$; in fact, one can show [23, Prop. 4.4] that $\lim_{\beta\rightarrow\infty}F_{\beta}(\boldsymbol{u},H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}))=\text{proj}_{T_{\mathcal{F}}(\boldsymbol{u})}\{-\nabla C_{p}(\boldsymbol{u})-J_{H}^{\top}\nabla C_{v}(H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}))\}$.
The algorithm (7) is designed to steer the power setpoints of the DERs $\boldsymbol{u}$ to optimal solutions of the AC OPF, while continuously guaranteeing feasibility (i.e., satisfaction of voltage limits). As shown in Figure 1, (7) effectively acts as a feedback controller by replacing the voltage model $H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$ with measurements $\tilde{\boldsymbol{\nu}}$ of the voltage magnitudes. This is a key modification that allows one to avoid collecting measurements of $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$ [7, 8, 14]. However, we note that (7) requires the computation of the Jacobian matrix of $H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$. When this is not possible, we modify it as follows:
@@
$\hskip 72.55461pt\nabla\hat{H}_{i}(\boldsymbol{u})^{\top}\boldsymbol{\theta}\leq-\beta\left(\tilde{\nu}_{i}-\bar{V}\right)\,\,\forall i\in\mathcal{M}$
@@
where $J_{\hat{H}}$ and $\{\hat{H}_{i}\}_{i\in\mathcal{M}}$ are computed based on a linear approximation of the AC power flow equations of the form
@@
$n\in\mathcal{N}$, where the coefficients $\{r_{n,m(i)},b_{n,m(i)}\}_{i\in\mathcal{G}}$ can be found as explained in e.g., [7, 26, 27, 32, 33] (note that linear approximations of the power flow equations are available for both single-phase and unbalanced multi-phase distribution networks). Effectively, the approximate Jacobian $J_{\hat{H}}$ no longer depends on $\boldsymbol{u}$ and the non-controllable powers $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$; accordingly, it does not need to be re-computed when running (9).
@@
Real-time operation: for $t\geq 0$
[S1a] Measure power setpoints $\{p_{i}(t),q_{i}(t),i\in\mathcal{G}\}$
@@
The safe gradient flow (9) relies on measurements of the voltages at the network locations $\mathcal{M}$. However, our framework is applicable to the case where the system operator may utilize a mix of actual voltage measurements and pseudo-measurements [34]. For example, suppose that the system operator can measure voltages at some nodes $\mathcal{M}_{\text{meter}}\subset\mathcal{M}$ and relies on pseudo-measurements at the other nodes. Then,

### v3

@@ Abstract
This paper considers the problem of controlling inverter-interfaced distributed energy resources (DERs) in a distribution grid to solve an AC optimal power flow (OPF) problem in real time. The AC OPF includes voltage constraints, and seeks to minimize costs associated with the economic operation, power losses, or the power curtailment from renewables. We develop an online feedback optimization method to drive the DERs’ power setpoints to solutions of an AC OPF problem based only on voltage measurements (and without requiring measurements of the power consumption of non-controllable assets). The proposed method – grounded on the theory of control barrier functions – is based on a continuous approximation of the projected gradient flow, appropriately modified to accommodate measurements from the power network. We provide results in terms of local exponential stability, and assess the robustness to errors in the measurements and in the system Jacobian matrix. We show that the proposed method ensures anytime satisfaction of the voltage constraints when no model and measurement errors are present; if these errors are present and are small, the voltage violation is practically negligible. We also discuss extensions of the framework to virtual power plant setups and to cases where constraints on power flows and currents must be enforced. Numerical experiments on a 93-bus distribution system and with realistic load and production profiles show a superior performance in terms of voltage regulation relative to existing methods.
Index Terms:
Distributed energy resources, AC optimal power flow, distribution networks, real-time control.
@@ I Introduction
Traditional techniques based on solving an AC optimal power flow (OPF) problem [3, 4] require collecting information of all non-controllable powers and running an iterative method; this process may be long compared to the fast changing conditions of a modern distribution system [1, 2]; existing Volt/Var techniques may not fully resolve voltage regulation and may in fact increase line currents; recent works on emulating OPF solutions via neural networks can alleviate the computational burden [5, 6], but still require measurements of all the non-controllable powers (which are the inputs to the neural network), and may not even produce feasible power setpoints.
Collecting measurements of non-controllable powers in distribution networks in real time (e.g., at the second level) is challenging as distribution systems are historically measurement-scarce, as discussed in [7]. Even though new types of sensing and communication infrastructures can capture real-time data, traditional techniques for solving the AC-OPF problem in distribution systems require collecting load measurements in real time from each meter and distribution transformer, which is impractical and economically unfeasible [8]. In this work, we focus on real-time AC OPF methods [9, 10, 11, 12, 13, 14], and seek new strategies that exhibit strong performance in terms of achieved operational cost and voltage limit satisfaction (both from analytical and numerical standpoints), while using limited measurements. In particular, we seek methods that do not require a complete AC model and knowledge of all the non-controllable powers throughout the nodes of the system.
@@
Contributions. Compared to the works in the context of real-time OPF methods mentioned above, the contributions of our paper can be described as follows:
@@
(c2) From a theoretical standpoint, we show that the proposed feedback-based SGF renders isolated optimal solutions of the AC OPF problem locally exponentially stable and ensures the anytime satisfaction of the voltage constraints. On the other hand, existing feedback-based optimization methods for distribution systems [9, 10, 11, 12, 13, 14] do not guarantee anytime satisfaction of the voltage constraints.
(c3) We provide results in terms of practical exponential stability and practical forward invariance when voltage measurements are affected by errors and when the Jacobian matrix of the AC power flow equations is computed only approximately (for example, when a linear approximation of the power flow equations is used).
(c4) We perform numerical experiments on a 93-bus distribution system [26] and with realistic load and solar production profiles from the Open Power System Data. We show that our method shows far superior performance in terms of voltage regulation relative to existing online primal-dual methods and Volt/Var strategies.
@@
From a practical standpoint, we provide remarks throughout the paper on how the proposed method can be integrated into the existing distribution system infrastructure: for instance, by leveraging distributed energy resources management systems (DERMSs) for distribution operators and existing communication infrastructure or a Supervisory Control and Data Acquisition (SCADA) system.
@@ II-A Distribution system model
Consider an electrical distribution system with $N+1$ nodes and hosting $G$ distributed energy resources (DERs); these may include inverter-interfaced photovoltaic systems, energy storage systems, variable-speed drives, and electric vehicles, or small-scale generators if any. The node 00 is taken to be the substation or the point of common coupling, while $\mathcal{N}:=\{1,...,N\}$ is the set of remaining nodes. We consider a steady-state model where voltages and currents are represented in the phasor domain. Accordingly, let $v_{k}=\nu_{k}e^{j\delta_{k}}\in\mathbb{C}$, $\nu_{k}:=|v_{k}|$, and $i_{k}=|i_{k}|e^{j\psi_{k}}\in\mathbb{C}$ the line-to-ground voltage and current injected at node $i$, respectively. Moreover, the voltage at node 00 is set to $v_{0}=V_{0}e^{j\delta_{0}}$ [27].
@@
where $\boldsymbol{s}=\boldsymbol{p}_{\text{net}}+\text{j}\boldsymbol{q}_{\text{net}}\in\mathbb{C}^{N}$, with $\boldsymbol{p}_{\text{net}}=[p_{\mathrm{net},1},...,p_{\mathrm{net},N}]^{\top}$ and $\boldsymbol{q}_{\text{net}}=[q_{\mathrm{net},1},...,q_{\mathrm{net},N}]^{\top}$ vectors collecting the net active and reactive power injections at nodes $\mathcal{N}$. Note that $\boldsymbol{p}_{\text{net}}$ and $\boldsymbol{q}_{\text{net}}$ account for both the powers (injected or consumed) of the DERs and the aggregate powers of the non-controllable loads that are connected to each of the nodes $\mathcal{N}$. In particular, let $\boldsymbol{p}_{l}:=[p_{l,1},\dots,p_{l,N}]^{\top}\in\mathcal{W}_{p},\boldsymbol{q}_{l}:=[q_{l,1},\dots,q_{l,N}]^{\top}\in\mathcal{W}_{q}$ with a compact sets $\mathcal{W}_{p}\subset\mathbb{R}^{N}$ and $\mathcal{W}_{q}\subset\mathbb{R}^{N}$, be vectors collecting the net active and reactive power consumed at the nodes by non-controllable devices (positive when the power is consumed). For the $G$ DERs, consider the vector $\boldsymbol{u}=[p_{1},p_{2},...,p_{G},q_{1},q_{2},...,q_{G}]^{\top}$ collecting their active and reactive powers (with a positive sign denoting generation). Moreover, let ${\mathcal{G}}:=\{1,...,G\}$ be the index for the DERs, and define a function $m:{\mathcal{G}}\rightarrow\mathcal{N}$ which maps a DER index to the node where the DER is connected to. With this notation, note that ${\mathcal{G}}_{n}:=\{i\in{\mathcal{G}}:n=m(i)\}$ is the set of DERs connected at node $n\in\mathcal{N}$. Then, the net active and reactive powers are given by $p_{\text{net},n}=\sum_{j\in{\mathcal{G}}_{n}}p_{j}-p_{l,n}$ and $q_{\text{net},n}=\sum_{j\in{\mathcal{G}}_{n}}q_{j}-q_{l,n}$ at each $n$. In what follows, we consider a set of nodes $\mathcal{M}\subseteq\mathcal{N}$ with cardinality $M=|\mathcal{M}|$ where voltages are to be regulated (if the operator would like to monitor and regulate all the voltages, then $\mathcal{M}=\mathcal{N}$). Finally, let $\mathcal{U}_{i}\subset\mathbb{R}^{2}$ be a compact set of admissible power setpoints, such that $(p_{i},q_{i})\in\mathcal{U}_{i}$ for $i\in\mathcal{G}$. We define $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}\times\ldots\times\mathcal{U}_{G}$ so that $\boldsymbol{u}\in\mathcal{U}$.
@@
There exists a unique continuously differentiable function $H:\mathcal{U}\times\mathcal{W}_{p}\times\mathcal{W}_{q}\rightarrow\mathbb{R}^{M}$ such that, $H_{i}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})=\nu_{i}=|v_{i}|$, for $i\in\mathcal{M}$. The Jacobian $J_{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}):=\frac{\partial H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})}{\partial\boldsymbol{u}}$ is locally Lipschitz continuous. $\Box$
@@
Remark II.1 (Jacobian of map $H$).
For the sake of generality, we write the Jacobian $J_{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$ as dependent on the controllable and non-controllable power injections. However, in the paper we will leverage approximations of the Jacobian matrix; these estimates can be obtained without any knowledge of non-controllable power injections. We will provide remarks on linear approximations in Section III-B. $\Box$
Remark II.2 (Model and notation).
@@
II-B OPF for voltage regulation in distribution systems
@@
To outline our framework, we start with the following formulation of the OPF problem (we present some extensions later in Section III-D):
@@
i) Strict complementarity slackness [31] and the linear independence constraint qualification (LICQ) [32] hold at $\boldsymbol{u}^{*}$.
@@
Problem (4) can be solved using traditional optimization methods for nonlinear programs. However, (c1) these batch methods require collecting measurements of all the non-controllable powers $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$ in a distribution network that is known to be historically measurement-scarce [7]; moreover, (c2) the time required to collect the measurements of the non-controllable powers (if available in real time) and run an iterative method to convergence may be long compared to the fast changing conditions of a modern distribution system [1, 2]. Although recent work on neural networks for AC OPF can alleviate the computational burden (see, e.g., the representative works [5, 6]), they still require measurements of all the non-controllable powers $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$ as in (c1), and they often rely on heuristics to return a feasible solution. Motivated by the challenges (c1)(c2) and by the need to generate power setpoints that ensure satisfaction of voltage limits even under uncertain and time-varying operational setups, in this paper we seek to solve to the following problem.
@@
Design an online feedback-optimization algorithm that drives the DERs’ power setpoints $\boldsymbol{u}$ to solutions of the problem (4), while ensuring that voltage constraints are always met. The feedback optimization method should use measurements of the voltages instead of requiring knowledge of the non-controllable powers $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$. $\Box$
@@
From a practical standpoint, in this paper we focus on feedback-optimization algorithms that are implemented in a centralized unit; for instance, these algorithms can be integrated into a DERMS for distribution operators. We also assume that the unit implementing our feedback-based online algorithm has access to synchronized voltage measurements at nodes where voltage constraints are enforced, and can transmit new power setpoints to the DERs. This can be done by leveraging existing communication and metering infrastructure or through a SCADA system.
@@ III-A Feedback-based online algorithm
$F_{\beta}(\boldsymbol{u},\tilde{\boldsymbol{\nu}}):=$
$\arg\min_{\boldsymbol{\theta}\in\mathbb{R}^{2G}}\|\boldsymbol{\theta}+\nabla C_{p}(\boldsymbol{u})+J_{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})^{\top}\nabla C_{v}(\tilde{\boldsymbol{\nu}})\|^{2}$
@@
$\hskip 42.67912pt\nabla H_{i}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})^{\top}\boldsymbol{\theta}\leq-\beta\left(\tilde{\nu}_{i}-\bar{V}\right)\quad\forall i\in\mathcal{M}$
$\hskip 42.67912ptJ_{\ell_{i}}(p_{i},q_{i})^{\top}\boldsymbol{\theta}\leq-\beta\ell_{i}(p_{i},q_{i})\qquad\qquad\forall i\in\mathcal{G}$
where $\tilde{\nu}_{i}$ is a measurement of $|v_{i}|$ at node $i$, $J_{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$ is the Jacobian matrix of $H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$, $\nabla H_{i}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})=[\{(J_{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}))_{i,j}\}_{j\in\mathcal{G}}]^{\top}$ is a $2G\times 1$ vector collecting the entries of $J_{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$ in the $i$th row and columns corresponding to nodes in $\mathcal{G}$, $J_{\ell_{i}}(p_{i},q_{i})$ is the Jacobian of $(p_{i},q_{i})\mapsto\ell_{i}(p_{i},q_{i})$, $\beta>0$ is a design parameter, and $\eta>0$ is the controller gain and is a design parameter. For given $\boldsymbol{u}$ and $\tilde{\boldsymbol{\nu}}$, problem (8) is a convex quadratic program (QP) with a strongly convex cost; it can be efficiently solved using standard or high-performance embedded solvers for QPs, e.g., [33].
The online feedback optimization algorithm (7) is inspired by CBFs methods [24] and the safe gradient flow in [25]; we provide more details on the CBF-based design in Appendix -A. In particular, (7) is an approximation of the projected gradient flow $\dot{\boldsymbol{u}}=\text{proj}_{T_{\mathcal{F}}(\boldsymbol{u})}\{-\nabla C_{p}(\boldsymbol{u})-J_{H}^{\top}\nabla C_{v}(H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}))\}$, where $T_{\mathcal{F}}(\boldsymbol{u})$ is the tangent cone of $\mathcal{F}(\boldsymbol{u})$ at $\boldsymbol{u}$; in fact, one can show [25, Prop. 4.4] that $\lim_{\beta\rightarrow\infty}F_{\beta}(\boldsymbol{u},H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}))=\text{proj}_{T_{\mathcal{F}}(\boldsymbol{u})}\{-\nabla C_{p}(\boldsymbol{u})-J_{H}^{\top}\nabla C_{v}(H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}))\}$.
The algorithm (7) is designed to steer the power setpoints of the DERs $\boldsymbol{u}$ to optimal solutions of the AC OPF, while continuously guaranteeing feasibility (i.e., satisfaction of voltage limits). As shown in Figure 1, (7) effectively acts as a feedback controller by replacing the voltage model $H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$ with measurements $\tilde{\boldsymbol{\nu}}$ of the voltage magnitudes that automatically satisfy the power flow equations [21, 34]. This is a key modification that allows one to avoid collecting measurements of $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$ [9, 10, 16]. However, we note that (7) requires the computation of the Jacobian matrix of $H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$. One cannot derive an explicit formulation of the Jacobian matrix of $H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})$ as this map does not have an analytical formulation. Therefore, we modify (7) as follows:
@@
$\hskip 72.55461pt\nabla\hat{H}_{i}^{\top}\boldsymbol{\theta}\leq-\beta\left(\tilde{\nu}_{i}-\bar{V}\right)\,\,\forall i\in\mathcal{M}$
@@
where $J_{\hat{H}}$ and $\{\hat{H}_{i}\}_{i\in\mathcal{M}}$ are estimates or approximations of $J_{H}$ and $\{H_{i}\}_{i\in\mathcal{M}}$, respectively. These estimates can be obtained using online estimation methods (see, e.g., [35, 36]), or they can be computed based on a linear approximation of the AC power flow equations of the form
@@
$n\in\mathcal{N}$, where the coefficients $\{r_{n,m(i)},b_{n,m(i)}\}_{i\in\mathcal{G}}$ can be found as explained in e.g., [9, 28, 29, 37, 38] (note that linear approximations of the power flow equations are available for both single-phase and unbalanced multi-phase distribution networks). We will provide more remarks on the linear approximation shortly in Section III-B.
Figure 1: Closed-loop implementation of the proposed online feedback optimization algorithm. A central unit (blue box) receives measurements of voltages (from measurement units, green box) and of the DERs’ output powers from inverters (red box); based on these measurements, it updates the DERs’ setpoints based on the proposed controller $\dot{\boldsymbol{u}}(t)=\eta\hat{F}_{\beta}(\boldsymbol{u}(t),\tilde{\boldsymbol{\nu}}(t))$. Once the setpoints $\boldsymbol{u}(t)$ are computed, the central unit transmits $\boldsymbol{u}(t)$ to the DERs’ inverters. Through this closed-loop scheme, the proposed controllers drive the distribution system to solutions of the AC OPF problem (4).
@@
Real-time operation: for $t\geq 0$, repeat:
[S1a] Measure output powers $\{p_{i}(t),q_{i}(t),i\in\mathcal{G}\}$
@@
Go to [S1a] and [S1b]
In terms of implementation of the Algorithm 1, we highlight the following practical aspects:
  • The main step [S2a] is performed at a central unit (i.e., the blue box in Fig. 1). This central unit can be integrated, for example, into a DERMS or an advanced distribution management system for distribution operators. After performing step [S2a], the central unit sends the updated setpoints to the DERs’ inverters.
  • Step [S1a] is performed at the DERs (a DER is represented by a red box in Fig. 1); the inverters measure the output powers $\{p_{i}(t),q_{i}(t),i\in\mathcal{G}\}$ and send the measurements to the central unit. The inverters also implement step [S2b] after they receive the setpoints from the central unit.
  • The safe gradient flow (9) relies on measurements of the voltages at the network locations $\mathcal{M}$, as required in the step [S1b]. It is assumed that those measurements are obtained in real time using sensing devices communicating with the centralized controller, e.g., $\mu$PMUs [39] or through the advanced metering infrastructure (a meter is represented by a green box in Fig. 1).
In practice, the proposed measurement-based SGF (9) can be implemented with discretization (similar to well-established CBF-based methods [24]). The discretization interval depends on the time required to collect voltage measurements and to solve the QP. Linearly-constrained convex QP programs are known to be solved efficiently (e.g, in milliseconds) by both existing open-source solvers (such as IPOPT) and commercial solvers. Synchronized voltage measurements can be obtained via SCADA at fast scale (i.e., second or sub-second level) [7, 39].
@@
Our framework is applicable to the case where the system operator may utilize a mix of actual voltage measurements and pseudo-measurements [40]. For example, suppose that the system operator can measure voltages at some nodes $\mathcal{M}_{\text{meter}}\subset\mathcal{M}$ and relies on pseudo-measurements at the other nodes. Then,